Calculate the work done by a system when it expands against a constant external pressure of 4.6 atm from an initial volume of 6.9 L to a final volume of 21.6 L.

Give your answer in joules (J) to the nearest integer (whole number). Do NOT input units. If your answer is negative, use a hyphen - (do not use an underscore) and do not add any spaces. For example, if your answer is -345.678 J, enter -346
Answer #1

Work done = - P(Vf - Vi)

External pressure, P = 4.6 atm

Final volume, Vf = 21.6 L

Initial volume, Vi = 6.9 L

Work done = - 4.6(21.6 - 6.9) L.atm

= - 67.62 L.atm

As, 1 L.atm = 101.3 J

Thus, work = -101.3×67.62 J

= - 6849.9 J

= - 6850 J